# Structured Generative Modeling with The Thermodynamic Kolmogorov-Arnold Model

### v4

@@ 1.1 Inductive biases and fast sampling
To address these limitations, we reimagine the EBM prior into a novel univariate or element-wise model, drawing inspiration from Kolmogorov-Arnold Networks (KANs) (Liu et al., 2024) and deep spline networks Bohra et al. (2020), which replace the sigmoidal activations used in multilayer perceptrons Murtagh (1991) with spline interpolations. By representing the latent space with univariate functions and distributions, we enable fast and exact prior sampling with the inverse transform method (Devroye, 1986) and improve interpretability for domain experts who may not be deeply familiar with computational statistics and machine learning. This allows the EBM prior to be aligned with the posterior before training begins.
@@ 1.2 Population-based Langevin sampling and Thermodynamic training
The resulting training criterion is derived from the discretized Thermodynamic Integral and the Steppingstone Estimator (SE), as presented by Calderhead & Girolami (2009); Annis et al. (2019) to ensure the impact of annealing on learning dynamics and parameter updates. Although this reduces expressivity compared to hierarchical EBMs, it ensures that inductive biases remain easy to embed into the latent prior and interpret after training; and unlike denoising, annealing is embarrassingly parallel.
@@ 2.1 Univariate energy-based prior
is measurable. Here, $\delta_{\psi_{q,p}(u_{p})}$ denotes the Dirac measure centred at $\psi_{q,p}(u_{p})$, ensuring the kernel is deterministic: i.e., once $u_{p}$ is sampled, the outcome in $\mathcal{Z}$ is fixed. This guarantees that the pushforward from $([0,1],\mathcal{B}([0,1]),\mu)$ to $(\mathcal{Z},\mathcal{B}(\mathcal{Z}),\pi)$ is well defined and that no additional stochasticity is introduced beyond the uniform source variable, $u_{p}$. Here, $\mathcal{B}(\mathcal{Z})$ is the Borel $\sigma$-algebra on the latent space, $\mathcal{Z}$, and $\pi$ is the latent measure.
The functions $\psi_{q,p}$ can therefore be conveniently interpreted as inverse cumulative distribution functions (CDFs) corresponding to independent univariate priors:
@@
Here, $F^{-1}(\pi_{q,p}):[0,1]\to\mathcal{Z}$ is the generalized inverse CDF associated with a parameterized latent distribution, $\pi_{q,p}$. This is valid under standard regularity assumptions on $F$, since it is a measurable, monotone, increasing function, and thus $F^{-1}(u)=\inf\left\{z\in\mathcal{Z}:F(z)\geq u\right\}$ is a measurable mapping between the two probability spaces. We parameterize the latent distribution $\pi_{q,p}$ via exponential tilting of a base prior $\pi_{0}$:
@@
where $f_{q,p}(z)$ is a learned energy function and $Z_{q,p}$ is the normalizing constant (partition function) that ensures $\pi_{q,p}$ defines a valid probability. The inspiration for this exponential tilting form can be traced to the energy-based prior introduced by Pang et al. (2020), adapted to the univariate case. The base prior $\pi_{0}$ serves as a reference measure that guides the form of the learned distribution $\pi_{q,p}$, allowing for flexible reweighting via the energy function $f_{q,p}$.
For $\pi_{q,p}$ to define a valid probability distribution, it must satisfy countable additivity while being non-negative and normalized. This is guaranteed since $f_{q,p}:\mathcal{Z}\to\mathbb{R}$ is Borel-measurable and integrable with respect to the reference density $\pi_{0}$, ensuring the finiteness of the partition function:
$Z_{q,p}=\int_{\mathcal{Z}}\exp(f_{q,p}(z))\,d\pi_{0}(z)<\infty.$ (5)
@@ 2.3.2 Mixture prior
As noted in Sec. 2.1, the independence of the priors $\pi_{q,p}$ limits expressivity. Moreover, the summation over $p$ in Eq. 8 accumulates the variances of sampled $z_{q,p}$, which can lead to overly diffuse representations. To address this, particularly for more complex generative tasks such as those in Sec. 3.2, we relax the strict adherence to KART by reformulating the prior as a mixture distribution. This involves moving the summation over $p$ out of the generator and absorbing it into the prior, yielding a mixture model:
@@
Here, $\alpha_{q,p}$ denote mixture weights, practically enforced via a softmax function. The resulting distribution defines a mixture of energy-based priors. Sampling becomes more efficient than in the previous formulation, as it proceeds component-wise using ITS for discrete variables, (Devroye, 2006). However, this violates adherence to KART, since $F$ now corresponds to a discrete rather than continuous CDF.
@@ 2.3.3 Deep KAN prior
If required, the energy functions $f_{q,p}$ can be deepened using the KAN approach. In fact, each $f_{q,p}$ can itself be a realization of Eq. 1 when its domain is restricted to $[0,1]$. This extension introduces new challenges:
@@ 2.4.3 Prior inverse transform sampling
In Eqs. 8 &  19, $z_{q,p}$ is drawn from $\pi_{q,p}$ with inverse transform sampling, (ITS) (Devroye, 1986). ITS can be used to sample from a component of the latent distribution, $\pi_{q,p}$, made possible given the unique characteristics of T-KAM. Specifically, the univariate, finite-grid structure of the constituent functions enables quadrature/interpolating integral approximations.
@@ 2.4.4 Posterior Expectations with Importance Sampling
The log-marginal likelihood and its learning gradient presented in Eqs. 15 &  16, involve expectations with respect to the posterior $P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi})$, which is potentially multimodal and complex. We propose Importance Sampling (IS) (Goertzel, 1949) as an efficient method for estimating its expectation when prior task knowledge can be incorporated into the model to improve prior-posterior alignment and the quality of the IS estimator.
@@
This is used to estimate the log-marginal likelihood and its gradient in Eqs. 15 and 16. However, unlike ULA (discussed in Sec. 2.5.2 and demonstrated in Sec. 3.2), IS fails to support posterior exploration when the prior is poorly aligned, making it unsuitable for general-purpose tasks beyond those similar to the datasets presented in Secs. 3.1.
@@ 3.2 Langevin Monte Carlo and Steppingstone estimator
We evaluated T-KAM using a CNN generator against two RGB image datasets at different resolutions, SVHN ($32\times 32$) (Netzer et al., 2011) and CelebA ($64\times 64$) (Liu et al., 2015). The first likelihood model in Eq. 20 was adopted for both datasets. Due to limited computational resources during this study, (detailed in Sec. A.2), training was shortened to 8,000 parameter updates, (70,000 were used in the study by Pang et al. (2020)), and smaller versions of the CNN decoders used by Pang et al. (2020) were adopted as the generator architectures.
Additionally, a single optimizer (Adam) was used to train both models, rather than using separate optimizers for the prior and generator networks. Chebyshev KAN bases (SS et al., 2024) were used for the prior energy functions, and full architectural details are provided under Sec. A.5.1.
Posterior sampling was conducted with ULA, and the standard normal reference prior was used, $\pi_{0}=\mathcal{N}\left(\bm{z};\;\bm{0},\bm{1}\right)$. T-KAM was trained for 40 epochs with a batch size of $N_{x}=100$ on 20,000 training examples. All remaining hyperparameters are provided in Tab. 6. Annealing was conducted with $N_{t}=10$ power posteriors and the cyclic temperature scheduling outlined in Sec. A.3.1.
@@
$\overline{\text{FID}}_{\infty}=97.90$ $\overline{\text{FID}}_{\infty}=145.64$
$\overline{\text{KID}}_{\infty}=0.0827$ $\overline{\text{KID}}_{\infty}=0.1407$
@@
$\overline{\text{FID}}_{\infty}=226.88$ $\overline{\text{FID}}_{\infty}=209.30$
$\overline{\text{KID}}_{\infty}=0.2590$ $\overline{\text{KID}}_{\infty}=0.2138$
@@
SE only resulted in better image quality for CelebA and cannot be justified in this setting given its computational overhead. Alternative approaches, such as different annealing schedules or the use of autoMALA with bias correction and temperature-adaptive step size tuning, could improve performance, but they are not worth investigating until the hardware of Zetta (2024) becomes available, due to the added complexity, hyperparameter tuning, and extended training time.
@@ 4.1.1 Hardware acceleration
Benchmarks are provided in Sec. A.2. The univariate, localized nature of spline interpolations complicates their parallelization on GPUs, since they involve sparse and recursive computations. The search operations in Eqs. 1323, and 69 were also difficult to parallelize on GPUs, which are generally ill-suited for search algorithms, as they often rely on conditional logic. Parallelizing conditionals can lead to thread divergence, making GPUs inefficient for these operations.
@@ 4.1.2 More opportunities for importance sampling
A possible remedy for the curse of dimensionality is to learn in a reduced-order space using Principal Component Analysis (PCA) (Jolliffe & Cadima, 2016), similar to the approach of Kingma et al. (2014), who also applied PCA to train VAEs. To address prior–posterior mismatch, future work might explore how existing domain knowledge can be incorporated to shape the prior’s energy functions and align them with the posterior before training begins. Additionally, efforts should be directed toward selecting an appropriate initial domain or adapting the domain dynamically during training with IS.
@@ A.2 Computational benchmarks
These benchmarks were collected for FP32 on a GPU (NVIDIA GeForce RTX 4060) with Julia’s BenchmarkTools package, (JuliaCI, 2024). The index search operations in Eqs. 2313, & 69 were parallelized across 20 threads on a CPU (i9-13900H @ 5.40GHz).
@@ A.3.1 Cyclic annealing
Instead, we schedule $p$ sinusoidally across all training parameter updates, starting with a higher value of $p$ and ending with a lower value, treated as tunable hyperparameters.
Figure 11: Cosine $p$ schedule, decreasing from 10 to 1
Initializing $p>1$ focuses learning on smoother likelihoods during the early stages of the cycle, where global features emerge between adjacent power posteriors. As training progresses, T-KAM will have learned the global features sufficiently well, reducing the change of measure between adjacent power posteriors of a $p>1$ schedule. At this stage, $p(i)$ will have transitioned to $p\leq 1$, thereby shifting T-KAM’s focus towards capturing more detailed features of the posterior landscape, where training is expected to enhance its ability to generate finer details.
@@ A.3.2 Choosing hyperparameters
$p_{\text{init}}>1,$ $p_{\text{end}}\leq 1,$ $N_{\text{cycles}}>0$ Controls the evolution of tempering during training. If the data is expected to have smooth relationships, reduce both $p_{\text{init}}$ and $p_{\text{end}}$. For data with finer details, increase both values to allow for better focus on capturing intricate features. Recommended values are $p_{\text{init}}=10,$ and $p_{\text{end}}=3$, since our experience suggests that power posteriors change rapidly closer to the prior. The number of cycles, $N_{\text{cycles}}$, controls the periodicity of cosine scheduling similar to cyclic annealing (Fu et al., 2019).
@@ A.5 Hyperparameters and training details
grid_range $z_{\text{min}},z_{\text{max}}]$ Uniform: $[0,1]$
@@
grid_range $z_{\text{min}},z_{\text{max}}]$ Uniform: $[0,1]$
@@
grid_range $z_{\text{min}},z_{\text{max}}] [$1,1][-1,1]
@@
GaussQuad_nodes $N_{\text{quad}}$ 200
@@
generator_variance $\sigma_{\epsilon}$ 0.25
@@
Number of epochs $N_{\text{epochs}}$ 40
@@
iters $N_{\text{local}}$ 40
step_size $\eta$ 0.002
@@
Initial exponent in power law schedule $p_{\text{start}}$ 10
Final exponent in power law schedule $p_{\text{end}}$ 3
@@ A.5.1 SVHN & CelebA Architectures
Input Sample $\bm{z}$ $p=20$
Energy functions Dense(2020) $q=41$

### v5

@@ 1.1 Inductive biases and fast sampling
To address these limitations, we reimagine the EBM prior into a novel univariate or element-wise model, drawing inspiration from Kolmogorov-Arnold Networks (KANs) (Liu et al., 2024) and deep spline networks Bohra et al. (2020), which replace the sigmoidal activations used in multilayer perceptrons Murtagh (1991) with spline interpolations. By representing the latent space with univariate functions and distributions, we enable fast, (single forward pass), and exact prior sampling with the inverse transform method (Devroye, 1986) and improve interpretability for domain experts who may not be deeply familiar with computational statistics and machine learning. This allows the EBM prior to be aligned with the posterior before training begins.
@@ 1.2 Population-based Langevin sampling and Thermodynamic training
The training criterion is derived from the discretized Thermodynamic Integral and the Steppingstone Estimator (SE), as presented by Calderhead & Girolami (2009); Annis et al. (2019) to ensure the impact of annealing on learning dynamics and parameter updates. Although this might be less expressive than hierarchical EBMs, it ensures that inductive biases remain easy to embed into the latent prior and interpret after training; and unlike denoising, annealing is embarrassingly parallel.
@@ 2.1 Univariate energy-based prior
is measurable. Here, $\delta_{\psi_{q,p}(u_{p})}$ denotes the Dirac measure centred at $\psi_{q,p}(u_{p})$, ensuring the kernel is deterministic: i.e., once $u_{p}$ is sampled, the outcome in $\mathcal{Z}$ is fixed. This ensures the pushforward,
$\left([0,1],\mathcal{B}\left([0,1]\right),\mu\right)\;\to\;\left(\mathcal{Z},\mathcal{B}\left(\mathcal{Z}\right),\pi\right),$ (3)
is well defined and that no additional stochasticity is introduced beyond the uniform source variable, $u_{p}$. Here, $\mathcal{B}(\mathcal{Z})$ is the Borel $\sigma$-algebra on the latent space, $\mathcal{Z}$, and $\pi$ is the latent measure. The functions $\psi_{q,p}$ can therefore be conveniently interpreted as inverse cumulative distribution functions (CDFs) corresponding to independent univariate priors:
@@
Here, $F^{-1}(\pi_{q,p}):[0,1]\to\mathcal{Z}$ is the generalized inverse CDF associated with a parameterized latent distribution, $\pi_{q,p}$. This is valid under standard regularity assumptions on $F$, since it is a measurable, monotone, increasing function, and thus $F^{-1}(u)=\inf\left\{z\in\mathcal{Z}:F(z)\geq u\right\}$ is a measurable mapping between the two probability spaces.
We parameterize the latent distribution $\pi_{q,p}$ via exponential tilting of a base prior $\pi_{0}$:
@@
where $f_{q,p}(z)$ is a learned energy function and $Z_{q,p}$ is the finite normalizing constant (partition function) that ensures $\pi_{q,p}$ defines a valid probability. The inspiration for this exponential tilting form can be traced to the energy-based prior introduced by Pang et al. (2020), adapted to the univariate case. The base prior $\pi_{0}$ serves as a reference measure that guides the form of the learned distribution $\pi_{q,p}$, allowing for flexible reweighting via the energy function $f_{q,p}$.
For $\pi_{q,p}$ to define a valid probability distribution, it must satisfy countable additivity while being non-negative and normalized. Assuming $f_{q,p}:\mathcal{Z}\to\mathbb{R}$ is $\mathcal{B}(\mathcal{Z})$-measurable with respect to the reference density $\pi_{0}$, this requires exponential integrability, i.e.,
$\exp(f_{q,p}(z))$ $\in L^{1}\left(\pi_{0}\right)$
$\text{such that}\quad Z_{q,p}$ $=\int_{\mathcal{Z}}\exp(f_{q,p}(z))\,d\pi_{0}(z)<\infty.$ (6)
@@ 2.3.2 Mixture prior
As noted in Sec. 2.1, the independence of the priors $\pi_{q,p}$ limits expressivity. Moreover, the summation over $p$ in Eq. 9 accumulates the variances of sampled $z_{q,p}$, which can lead to overly diffuse representations. To address this, particularly for more complex generative tasks such as those in Sec. 3.2, we relax the strict adherence to KART by reformulating the prior as a mixture distribution, similar to the approach described by Rivera (2024). This involves moving the summation over $p$ out of the generator and absorbing it into the prior, yielding a mixture model:
@@
Here, $\alpha_{q,p}$ represent the mixture proportions, which must satisfy non-negativity and unit-sum constraints. In this study, they were fixed uniformly as $1/n_{z}$, guided by Rivera (2024). They could alternatively be learned by passing trainable parameters through a softmax function, (to enforce the constraints), however this approach was not used for the experiments in Sec. 3.
The resulting distribution defines a mixture of energy-based priors. Sampling becomes more efficient than in the previous formulation, as it proceeds component-wise using ITS for discrete variables, (Devroye, 2006). However, this violates adherence to KART, since $F$ now corresponds to a discrete rather than continuous CDF.
@@ 2.3.3 Deep KAN prior
If required, the energy functions $f_{q,p}$ can be deepened using the KAN approach instead of the mixture approach in Sec. 2.3.2. In fact, each $f_{q,p}$ can itself be a realization of Eq. 1 when its domain is restricted to $[0,1]$. This would introduce new challenges:
@@ 2.4.3 Prior inverse transform sampling
In Eqs. 9 &  20, $z_{q,p}$ is drawn from $\pi_{q,p}$ with inverse transform sampling, (ITS) (Devroye, 1986). ITS can be used to sample from a component of the latent distribution, $\pi_{q,p}$, made possible given the unique characteristics of T-KAM. Specifically, the univariate, finite-grid structure of the constituent functions enables quadrature/interpolating integral approximations.
@@ 2.4.4 Posterior Expectations with Importance Sampling
The log-marginal likelihood and its learning gradient presented in Eqs. 16 &  17, involve expectations with respect to the posterior $P(\bm{z}\mid\bm{x}^{(b)},\bm{f},\bm{\Phi})$, which is potentially multimodal and complex. We propose Importance Sampling (IS) (Goertzel, 1949) as an efficient method for estimating its expectation when prior task knowledge can be incorporated into the model to improve prior-posterior alignment and the quality of the IS estimator.
@@
This is used to estimate the log-marginal likelihood and its gradient in Eqs. 16 and 17. However, unlike ULA (discussed in Sec. 2.5.2 and demonstrated in Sec. 3.2), IS fails to support posterior exploration when the prior is poorly aligned, making it unsuitable for general-purpose tasks beyond those similar to the datasets presented in Secs. 3.1.
@@ 3.2 Langevin Monte Carlo and Steppingstone estimator
We evaluated T-KAM using a CNN generator against two RGB image datasets at different resolutions, SVHN ($32\times 32$) (Netzer et al., 2011) and CelebA ($64\times 64$) (Liu et al., 2015). The first likelihood model in Eq. 21 was adopted for both datasets. Due to limited computational resources during this study, (detailed in Sec. A.2), training was shortened to 4,000 parameter updates, (70,000 were used in the study by Pang et al. (2020)), and smaller versions of the CNN decoders used by Pang et al. (2020) were adopted as the generator architectures.
A single optimizer (Adam) was used to train both models, rather than using separate optimizers for the prior and generator. FFT bases were used for the prior energy functions, and full architectural details are provided under Sec. A.5.1. Posterior sampling was conducted with 20 steps of ULA, and no reference prior was used, $\pi_{0}=1$. T-KAM was trained for 20 epochs with a batch size of $N_{x}=100$ on 20,000 training examples. All remaining hyperparameters are in Tab. 6. Annealing was conducted with $N_{t}=10$ power posteriors and the $p$ scheduling outlined in Sec. A.3.1.
@@
$\overline{\text{FID}}_{\infty}=120.53$ $\overline{\text{FID}}_{\infty}=167.79$
$\overline{\text{KID}}_{\infty}=0.1103$ $\overline{\text{KID}}_{\infty}=0.1644$
@@
$\overline{\text{FID}}_{\infty}=174.69$ $\overline{\text{FID}}_{\infty}=173.78$
$\overline{\text{KID}}_{\infty}=0.1917$ $\overline{\text{KID}}_{\infty}=0.1851$
@@
Our proposed SE training strategy resulted in slightly better image metrics for CelebA, but significantly worsened them with SVHN. However, from our subjective perspective, the generated images appeared more well-defined, yet less diverse under SE than under MLE. Ultimately, we cannot justify the Thermodynamic training strategy due to its substantial computational overhead. With more available VRAM than in this study, it would be possible to parallelize a greater number of temperatures than $N_{t}=10$, which could potentially improve quality, as suggested by the results of Calderhead & Girolami (2009).
More posterior sampling options, such as modified annealing schedules, a larger number of ULA steps, or the use of autoMALA with bias correction and temperature-adaptive step size tuning may further improve results. Nevertheless, exploring these options is not worthwhile until the hardware described by Zetta (2024) becomes available, given the added algorithmic complexity, hyperparameter sensitivity, and extended training times involved.
@@ 4.1.1 Hardware acceleration
Benchmarks are provided in Sec. A.2. The univariate, localized nature of spline interpolations complicates their parallelization on GPUs, since they involve sparse and recursive computations. The search operations in Eqs. 1424, and 70 were also difficult to parallelize on GPUs, which are generally ill-suited for search algorithms, as they often rely on conditional logic. Parallelizing conditionals can lead to thread divergence, making GPUs inefficient for these operations.
@@ 4.1.2 More opportunities for importance sampling
A possible remedy for the curse of dimensionality is to learn in a reduced-order space using Principal Component Analysis (PCA) (Jolliffe & Cadima, 2016), similar to the approach of Kingma et al. (2014), who also applied PCA to train VAEs. To address prior–posterior mismatch, future work might explore how existing domain knowledge can be incorporated to shape the prior’s energy functions and align them with the posterior before training begins. Additionally, efforts should be directed toward selecting appropriate initial domains or adapting the domains dynamically.
@@ A.2 Computational benchmarks
These benchmarks were collected for FP32 on a GPU (NVIDIA GeForce RTX 4060) with Julia’s BenchmarkTools package, (JuliaCI, 2024). The index search operations in Eqs. 2414, & 70 were parallelized across 20 threads on a CPU (i9-13900H @ 5.40GHz).
@@ A.3.1 Cyclic annealing
Instead, we schedule $p$ sinusoidally across all training parameter updates, starting with a lower value of $p$ and ending with a higher value, treated as tunable hyperparameters.
Figure 11: Cosine $p$ schedule, increasing from 1 to 8
Initializing $p\leq 1$ focuses learning on high-variance, data-informed likelihoods during the early stages of the cycle. Here, features emerge closer to the true posterior, and learning requires exploration and higher variance gradient estimates. As training progresses, T-KAM will have learned global features sufficiently well, reducing the meaningful change of measure between adjacent power posteriors of a $p\leq 1$ schedule. At this stage, $p(i)$ will have transitioned to $p>1$, thereby shifting focus towards exploitation and sculpting the prior with lower variance gradient estimates.
@@ A.3.2 Choosing hyperparameters
$p_{\text{init}}\leq 1,$ $p_{\text{end}}>1,$ $N_{\text{cycles}}\geq 0$ Controls the evolution of tempering during training. If the data is expected to have smooth relationships, reduce both $p_{\text{init}}$ and $p_{\text{end}}$. For data with finer details, increase both values to allow for better focus on capturing intricate features. The values used for this study were $p_{\text{init}}=1$, and $p_{\text{end}}=8$, since our experience suggests that power posteriors change rapidly closer to the prior. The number of cycles, $N_{\text{cycles}}$, controls the periodicity of cosine scheduling similar to cyclic annealing (Fu et al., 2019).
@@ A.5 Hyperparameters and training details
Initial grid_range (before grid updating) $z_{\text{min}},z_{\text{max}}]$ Uniform: $[0,1]$
@@
Initial grid_range (before grid updating) $z_{\text{min}},z_{\text{max}}]$ Uniform: $[0,1]$
@@
Initial grid_range (before grid updating) $z_{\text{min}},z_{\text{max}}] [$1.5,1.5][-1.5,1.5]
@@
GaussQuad_nodes $N_{\text{quad}}$ 400
@@
generator_variance $\sigma_{\epsilon}$ 0.3
@@
Number of epochs $N_{\text{epochs}}$ 20
@@
iters $N_{\text{local}}$ 20
step_size $\eta$ 0.01
@@
Initial exponent in power law schedule $p_{\text{start}}$ 1
Final exponent in power law schedule $p_{\text{end}}$ 8
@@ A.5.1 SVHN & CelebA Architectures
Input Sample $\bm{z}$ $p=40$
Energy functions Dense($40$) $q=81$